If (r+ (1/r))^2=5, then find the value of r^3 + 1/r^3.
Answer by venugopalramana(3286)

(R+1/R)^2=5
(R+1/R)=5^0.5...CUBE BOTH SIDES
(R+1/R)^3=(5^0.5)^3
R^3+1/R^3 +3*R*(1/R)*(R+1/R)=5*SQRT5
R^3+1/R^3=5*SQRT(5)-3*SQRT(5)=2*SQRT(5)